OpenStudy (liizzyliizz):

4e^2x=12 (for this equation i got x=.5493 ) am i correct?

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

or in ln form x= ln3/2

myininaya (myininaya):

\[4e^{2x}=12=>\frac{4}{4}e^{2x}=\frac{12}{4}=>e^{2x}=3\] \[\ln(e^{2x})=\ln(3) =>2xln(e)=\ln(3)=>2x(1)=\ln(3)=>2x=\ln(3)=>x=\frac{\ln(3)}{2}\]